# Calibration Attack: A Framework For Adversarial Attacks Targeting Calibration

### v1

@@ 3.2 Four Types of Calibration Attack
Random Confidence Attack. This variation of attacks attempt to completely decouple the model’s confidence scores and predictive performance in a randomized fashion by modifying inputs, such that the confidence scores produced by the model on them are randomized. Specifically, random confidence attacks are performed by choosing a random goal confidence score for each input, and then, depending on the original confidence score, running the corresponding underconfidence or overconfidence attacks to produce a new target confidence score on the input. Although this form of attacks are theoretically less effective than the maximum miscalibration attacks, they are less predictable, since unlike the other attacks it does not lead the model to produce one or two terminal confidence scores for each input, and instead can produce more initially reasonable looking confidence scores that are completely meaningless due to being randomized.
3.3 Defence Against Calibration Attack
As we will show in Section 4, calibration attacks are very effective on existing victim models, hence developing defence strategies is important. In addition to investigating existing defence approaches against regular adversarial attacks, we propose two novel methods specifically against calibration attacks: Calibration Attack Adversarial Training (CAAT) and Compression Scaling (CS).
@@ 3.3 Defence Against Calibration Attack
Table 1: Results of four types of $l_{\infty}$ calibration attacks: underconfidence (und-atk), overconficence (ovr-atk), maximum miscalibration (max-atk), and random confidence attack (rnd-atk). Accuracies of victim models are included.
@@
Und-atk $169.8_{\pm 15.0}$ $88.3_{\pm 6.7}$$459_{\pm.015}.$$452_{\pm.019}.$$498_{\pm.026}.$
Ovr-atk $94.9_{\pm 45.9}$ $3.7_{\pm 4.6}$$029_{\pm.003}.$$030_{\pm.004}.$$976_{\pm.011}.$
Max-atk $137.1_{\pm 4.3}$ $88.3_{\pm 6.7}$$519_{\pm.020}.$$480_{\pm.020}.$$509_{\pm.024}.$
Rnd-atk $129.5_{\pm 7.4}$ $97.2_{\pm 9.9}$$454_{\pm.012}.$$432_{\pm.016}.$$538_{\pm.019}.$
CAAT is a variation of PGD-based adversarial training utilizing our white-box calibration attacks to generate adversarial training samples for each minibatch during training. Hence, both under and overconfident samples with the model’s original predicted label preserved are exclusively used to train the model. CS is a post-process scaling technique primarily designed for dealing with maximum miscalibration attacks. It is based on the assumption that since effective classifiers have a high level of accuracy (and confidence), calibration attacks will typically do most of the damage by lowering confidence scores on highly confident correctly classified samples. Therefore, by scaling the low scores to a high confidence range, CS can help mitigate the most severe miscalibration.
For our algorithm, the range of possible confidence scores are first split into $\bm{M}$ equally sized bins, which we select to be 3 (or 4). Any samples whose confidence scores fall within the bin $\bm{m}\in\{1,...,\bm{M}\}$ in original group $\bm{O}$ are mapped to a corresponding bin in high confidence bin group $\bm{Q}$ with a compressed confidence range. In our case, we divide confidence scores into 15 bins and chose the top 3 (or 4) highest confidence bins as the corresponding compressed bins. A linear mapping is applied and the samples’ positions in the new bin are kept to be the same as in the original one. Next, the sample’s logit vector is scaled by finding a scaling factor by iterating through a large range of possible values so that the new desired confidence score for the sample achieved within the new confidence range. Formally, for a sample $\bm{x}$ with logit distribution $l=\{l_{1},...,l_{k}\}$ and original predicted probability $\bm{\hat{p}}_{og}$, we find temperature $T$ such that it is mapped from a relative bin in bin group $\bm{O}$ to bin group $\bm{Q}$. The new predicted probability ${\hat{p}}_{new}$ is ${\operatorname*{arg\,max}}_{i}(\frac{{\exp}(l_{i}/T)}{{\sum}_{j}{exp}(l_{j})/T})$ such that ${\hat{p}}_{new}=\mathrm{min\>conf}({Q_{m}})+\frac{{\hat{p}}_{og}-\mathrm{min\>conf}({O_{m}})}{\mathrm{range}(O_{m})}*\mathrm{range}(Q_{m}),$ where $\mathrm{min\>conf}(Q_{m})$ is the minimum confidence level of bin $\bm{m}$ in $\bm{Q}$, while $\mathrm{max\>conf}(O_{m})$ is the maximum confidence level of bin $\bm{m}$ in $\bm{O}$, and $\mathrm{range}()$ gives the range of confidence vales in a bin. That is, for bin group $O$ it is $\mathrm{max\>conf}(O_{m})-\mathrm{min\>conf}(O_{m})$.
4 Experimental Results of Calibration Attack
Metrics. Two widely used metrics are used to assess calibration error: Expected Calibration Error (ECE) (Pakdaman Naeini et al., 2015) and Kolmogorov-Smirnov Calibration Error (KS error) (Gupta et al., 2021). We evaluate attacks’ query efficiency using the average and median number of queries for the attacks to complete (Andriushchenko et al., 2020). Average confidence of predictions is used to judge the degree that the confidence scores are affected. (See Appendix B for details.)
Datasets. We use CIFAR-100 (Krizhevsky & Hinton, 2009), Caltech-101 (Fei-Fei et al., 2004), and the German Traffic Sign Recognition Benchmark (GTSRB) (Houben et al., 2013) with respect to safety critical applications.
Models. To explore the effect of attacks on different architecture, we use ResNet-50 (He et al., 2016) and the popular non-convolutional attention-based model, the Vision Transformer (ViT) (Dosovitskiy et al., 2021). Details can be seen in Appendices B and C.
Attack Settings. Key attack settings, e.g., $l_{\infty}$, $l_{2}$ and attack iterations, can be found in Appendix B.
Figure 2: The influence of $\epsilon$ (the left three subfigures) and attack iterations (the right subfigure) in the attacks effectiveness. Subfigure-1 (the left most) presents the comparison between the ECE scores of the different calibration attacks at different $\epsilon$ values using ResNet-50 models trained on CIFAR-100. Subfigure-2: ECE vs. $\epsilon$ using maximum miscalibration attacks on ViT models trained on CIFAR-100. Subfigure-3: ECE vs. $\epsilon$ using maximum miscalibration attacks on the ResNet-50 models trained on Caltech-101 and GTSRB. Subfigure-4: Effect of the numbers of attack iterations on the ability of the attack algorithm. The first three subfigures are created at the $1000^{th}$ iteration.
@@ 4.1 Overall Performance
The performance of $l_{\infty}$ attacks on different victim models are shown in Table 1. The results of $l_{2}$ attacks can be found in Table 6 of the Appendix F.6, and the main results for the white-box attacks are in Table 5 of the Appendix F.1. The experiments show that the maximum miscalibration attack is consistently the most effective in producing the highest levels of miscalibration in both black-box and white-box setups. In most cases its efficiency is competitive, suggesting it has overall the best performance under the same number of attack iterations. The underconfidence and random confidence attack also create substantial calibration error. They perform similarly to one another in terms of overall error, with the underconfidence attacks largely edging in in effectiveness. The results are obtained on three runs of each model with different random seeds.
The overconfidence attack results are of particular interest—the attack successfully raises the average confidence level to around 99% in most cases. Given the very high level of accuracy of each of these models, increasing the confidence level of the predictions does not have as drastic an effect on the calibration error since the gap between the average confidence level and accuracy will still be small when confidence levels are raised to 100%. It can even appear to make the model produce better calibrated scores on the new adversarial data if the model was originally underconfident on the data. However, the effect of the overconfidence attacks is still dangerous since it can make the model output near certain confidence scores irrespective of the input. In practice this can render the use of confidence scores meaningless as they would be the same for every input, so the low calibration error does not necessarily reflect the severity of the attacks. Figure 1 demonstrates the effects of each form of attacks in calibration diagrams, in which we can observe the severity of miscalibration.
The general trends between the $l_{2}$ and $l_{\infty}$ attacks are similar, although the latter are more effective, leading to more severe miscalibration whilst being more query efficient and faster by as much as two or three times. There are only a few cases where $l_{2}$ attacks yield higher miscalibration or more efficient attacks in terms of the number of queries. We focus our remaining analyses on the more effective $l_{\infty}$ norm. Regarding different architectures, the ViT models tend to be slightly more miscalibrated after the attacks, compared to the ResNet models on ECE and KS, but the ViT models also are notably more miscalibrated before the attacks.
Even with the relatively low amount of queries under the black-box setting, all of the types of attacks are generating severe miscalibration, increasing ECE and KS error often by over 10 times the original amount. Calibration attacks are very effective in achieving their goals even with the prediction accuracy unchanged, raising serious concerns for safety critical applications.
@@
Table 2: Adversarial attack detection results comparing original version of PGD and SA attacks with their calibration attack counterparts.
PGD PGD-Under PGD-Over PGD-Max SA SA-Under SA-Over SA-Max
@@
LID 95.7 89.1 88.8 86.3 94.6 87.4 87.0 85.3 86.3 77.1 71.4 68.1 72.5 69.8 72.4 66.7
MD 100.0 99.8 94.6 93.6 97.9 97.7 92.9 92.6 95.5 89.6 83.7 77.8 74.4 74.9 85.6 79.0
Spect. 99.4 98.5 94.8 93.0 99.0 99.9 99.0 99.9 99.1 98.0 83.0 79.0 50.4 50.5 83.1 79.0
@@
4.2 Detection Difficulty Analysis
Calibration attacks are difficult to detect given they do not affect the decisions of the classifier. To show this, we run experiments using ResNet-50 on our three tested datasets using the popular adversarial attack detection methods: Local Intrinsic Dimensionality (LID) (Ma et al., 2018), Mahalanobis Distance (MD) (Lee et al., 2018), and SpectralDefense (Harder et al., 2021). The details behind the settings for each detection method can be seen in Appendx E. We show our main results of the effectiveness of the detectors in terms of Area Under the Curve (AUC) and detection accuracy when running the different types of calibration attacks compared to their original counterparts in Table 2, under both the white-box and black-box setups. We observe how in most cases, with one exception, there is a notable decrease in detection performance, with the effects being most pronounced in SA, where the decrease of more than 20% in accuracy can often be seen. Existing detection methods appear less reliable for calibration attacks.
4.3 Impact of Key Aspects of Attack
We examine the impact of key aspects of calibration attacks, including adversarial noise, attack iterations, and attack types. More details can be found in Appendix F, including additional analyses. All the analyses are conducted over the more effective $l_{\infty}$ attacks in the black-box setup using our regular attack settings unless stated otherwise.
Epsilon. We study how successful the different attacks are at different $\epsilon$ levels to ascertain how low the value can be in order to have notable harm. The results are shown in Figure 2. Our first set of experiments seen in the leftmost figure is based on CIFAR-100 using ResNet-50, where we track the ECE at five different $\epsilon$ values after being attacked using all the different attacks. The attacks are successful even with low $\bm{\epsilon}$ values. The rise in ECE is sharp when $\bm{\epsilon}$ increases, but it plateaus at higher values. The middle figure shows similar results when using ViT with the maximum miscalibration attack, revealing that the phenomena is not architecture dependent. The rightmost figure, where a maximum miscalibration attack is used against ResNet over both Caltech-101 and GTSRB, shows similar patterns across different datasets.

### v2

@@ 3.2 Four Forms of Calibration Attacks
Random Confidence Attacks (RCA). We propose to perform RCA to decouple the victim model’s confidence scores and predictive labels, in which the confidence scores produced by the attack model are randomized. RCA is performed by choosing a random target confidence score for each input, and then, depending on the original confidence score, running the corresponding underconfidence or overconfidence attacks to produce the target confidence score. Although RCA is theoretically less effective than MMA, it is less predictable, because unlike the other three types of attacks, RCA does not produce adetermined direction (i.e., under or overconfidence) of attacks for a given input. The produced confidence scores are often less extreme and more reasonable-looking, but largely meaningless due to randomization. Note that in addition to MMA and RCA, one may design other approaches to combine UCA with OCA, but MMA and RCA represent the most typical composition.
3.3 Defence Against Calibration Attacks
As we will show in the experiments, calibration attacks are very effective. In addition to studying a wide range of existing defence methods, we introduce new methods specifically for calibration attacks: Calibration Attack Adversarial Training (CAAT) and Compression Scaling (CS). The proposed CAAT is a variation of PGD-based adversarial training that utilizes our white-box calibration attacks to generate adversarial training examples for each minibatch during training. Hence, both under- and overconfident examples with the model’s original predicted label preserved are exclusively used to train the model.
Our proposed CS defence is a post-process scaling technique, based on the assumption that an effective classifier often has a high level of accuracy and confidence, so calibration attacks typically cause the most harm by lowering the confidence scores. CS hence aims to scale low confidence scores to high values to mitigate the damage of miscalibration. Specifically, the range of confidence is split into $B_{M}$ equally sized bins. Datapoints in each bin $b_{m}\in\{1,...,B_{M}\}$ are mapped to a new bin that has higher confidences in a more compressed range. For a datapoint $\mathbf{x}_{i}$ with the output-layer logits $\{r_{1},...,r_{K}\}$ and the probability ${\hat{p}}_{k}(\mathbf{x}_{i})$ for the predicated class $k$, a temperature $T$ is found to obtain a newly predicted probability ${\hat{p}}_{new}={\operatorname*{arg\,max}}_{i}(\frac{{\exp}(r_{i}/T)}{{\sum}_{j}{exp}(r_{j})/T})$ so that ${\hat{p}}_{new}=\mathrm{minconf}(b^{\prime}_{m})+\frac{{\hat{p}}_{k}-\mathrm{minconf}({b_{m}})}{\mathrm{range}(b_{m})}*\mathrm{range}({b^{\prime}_{m}})$, where $\mathrm{minconf}(.)$ is the minimum confidence level of a bin, and $\mathrm{range}(.)$ represents the range of the confidence (i.e., the maximum level of confidence of that bin minus the minimum). $b^{\prime}$ is the bin that the original $b$ is mapped to using the corresponding $T$. We will demonstrate that even with defence, calibration attacks are still highly effective.
3.4 Discussion on the Importance of Remaining Well Calibrated Under the Attacks
In addition to the discussion of the technical details of calibration attacks for the victim models and defences, it is important to reemphasize why maintaining good calibration under such attacks is critical. Well-calibrated models are a crucial component for trustworthy complex systems. As an example, in real-life deployment, proper confidence scores are essential for determining the instances that need further examination by authority mechanisms such as an expert or a committee of them. Underconfidence attacks could cause additional strain on an authority system or any downstream processes, creating a risk of significantly slowed-down decision-making due to an increased load of cases to be examined. As another example, by attacking misclassified datapoints with an overconfidence approach, test cases may be erroneously missed by downstream systems such as braking in an autonomous vehicle before a stop sign.
4 Experiments
4.1 Experimental Setup
Implementation Deails. As discussed earlier in this paper, calibration attacks can be built on different adversarial attack frameworks. In our implementation we use Square Attack (SA)  (Andriushchenko et al., 2020), which is one of the most popular black-box approaches and is highly effective. SA achieves state-of-the-art (SOTA) performance in terms of query efficiency and success rate, even outperforming some white-box methods. Our white-box calibration attacks are based on the popular Projected Gradient Descent (PGD) framework (Madry et al., 2018). The implementation details such as hyperparameters are discussed in Appendix C.
Models. We include both convolutional (ResNet (He et al., 2016)) and non-convolutional attention-based models (Vision Transformer (ViT) (Dosovitskiy et al., 2021)) in our study. Details can be found in Appendices C and D.
Datasets. We performed a comprehensive study on CIFAR-100 (Krizhevsky & Hinton, 2009) and Caltech-101 (Fei-Fei et al., 2004). We also included the German Traffic Sign Recognition Benchmark (GTSRB) (Houben et al., 2013) for its direct implication for safety.
@@ 4.1 Experimental Setup
Table 1: Results of underconfidence, overconfidence, maximum miscalibration, and random confidence attack. Accuracy of victim models are included.
@@
UCA $169.8_{\pm 15.0}$ $88.3_{\pm 6.7}$$459_{\pm.015}.$$452_{\pm.019}.$$498_{\pm.026}.$
OCA $94.9_{\pm 45.9}$ $3.7_{\pm 4.6}$$029_{\pm.003}.$$030_{\pm.004}.$$976_{\pm.011}.$
MMA $137.1_{\pm 4.3}$ $88.3_{\pm 6.7}$$519_{\pm.020}.$$480_{\pm.020}.$$509_{\pm.024}.$
RCA $129.5_{\pm 7.4}$ $97.2_{\pm 9.9}$$454_{\pm.012}.$$432_{\pm.016}.$$538_{\pm.019}.$
Metrics. Two calibration metrics are used in our experiments: the stardard Expected Calibration Error (ECE) (Pakdaman Naeini et al., 2015) and the recent Kolmogorov-Smirnov Calibration Error (KS error) (Gupta et al., 2021). In addition, we evaluate attacks’ efficiency using the average and median number of queries for the attack to complete (Andriushchenko et al., 2020). Average confidence of predictions is also leveraged to judge the degree that the confidence scores are affected. (See Appendix C for details.)
Detailed Attack Settings. Descriptions of other attack settings (e.g., $l_{\infty}$, $l_{2}$ and iterations) are in Appendix Cand D.
@@ 4.2 Overall Performance
Table 1 depicts the overall performance of black-box attacks under the $l_{\infty}$ norm. Additional results of $l_{2}$ attacks are included in Table 6 in Appendix G.9. The experiments show that the attacks are highly effective. UCA, MMA and RCA can bring ECE and KS to very high values. OCA can successfully raise average confidences (the last column in Table 1) to extremely high levels (in many cases, close to 100%), but given the original high accuracy of the victim models, increasing confidence levels will not have a drastic effect on calibration error. (OCA will be more harmful on less accurate models.) For the MMA, the theoretical highest levels of miscalibration are not reached due to the limited number of iterations that we run the attacks. Regarding different architectures, the attention-based ViT models are seen to be more miscalibrated compared to the convolution-based ResNet models. The white-box results are included in Appendix G.1, which show a similar trend. Again, Figure 1 in the introduction section demonstrates the attack effects using calibration diagrams, visually showing the severity of miscalibration.
The general trend of the $l_{2}$ calibration attacks (Appendix G.9) is similar to that of the $l_{\infty}$ attacks, but the latter are found to be more effective in generating more significant miscalibration. Hence in the remainder of the paper, we focus on the $l_{\infty}$ attacks unless otherwise specified.
Overall, calibration attacks are generating severe miscalibration, which, compared to the pre-attack values, can increase ECE and KS by over 10 times in many cases. Calibration attacks are very effective without changing the prediction accuracy, which could raise serious concerns for any down-stream applications relying on confidence.
4.3 Detection Difficulty Analysis
This section shows calibration attacks are also difficult to detect. To investigate this, we run the popular adversarial attack detection methods on the attacks against ResNet-50: Local Intrinsic Dimensionality (LID) (Ma et al., 2018), Mahalanobis Distance (MD) (Lee et al., 2018), and SpectralDefense (Harder et al., 2021). The details behind the settings for each detection method can be found in Appendix F. Table 2 depicts the main results of the effectiveness of the detectors in terms of Area Under the Curve (AUC) and Detection Accuracy, under different types of calibration attacks and using both the white-box and black-box approaches. We can see that there are consistent decreases in detection performances, particularly in SA, where the decreases are often more than 20%. The existing detection methods are shown to be less reliable for calibration attacks.
4.4 Insights on Key Aspects of Attacks
We analyze key aspects of calibration attacks: attack directions, noise bounds, and attack iterations under the $l_{\infty}$-based attacks. More details can be found in Appendix G, including attack effectiveness on data with varying imbalance ratios.
@@ 4.4 Insights on Key Aspects of Attacks
Table 2: Attack detection results comparing original version of SA and PGD attacks with their calibration attack counterparts.
SA SA-UCA SA-OCA SA-MMA PGD PGD-UCA PGD-OCA PGD-MMA
@@
LID 86.3 77.1 71.4 68.1 72.5 69.8 72.4 66.7 95.7 89.1 88.8 86.3 94.6 87.4 87.0 85.3
MD 95.5 89.6 83.7 77.8 74.4 74.9 85.6 79.0 100 99.8 94.6 93.6 97.9 97.7 92.9 92.6
Spect. 99.1 98.0 83.0 79.0 50.4 50.5 83.1 79.0 99.4 98.5 94.8 93.0 99.0 99.9 99.0 99.9
[…]
Comparison of Efficiency of Underconfidence vs. Overconfidence Attacks. The two directions of calibration attacks, underconfidence or overconfidence, is a basic building block for constructing the four forms of calibration attacks. We perform further studies to understand which direction is most efficient. To the end, we identify all data points in the test set that are around certain predefined base confidence levels. In this study, we choose two base confidences: 80% and 90%. All the test cases that have confidences within 1% around these two base confidences are included in this experiment. We attack these examples by making the victim models produce either a 10% increase or decrease in confidence. In Table 3 we can see a consistent pattern — for both base confidence levels, it takes notably fewer queries to create underconfidence than overconfidence adversarial examples. The former attack is also more effective in affecting the average confidences. This property could be further utilized to design calibration attacks (e.g., under the circumstances where computing resources or attack latency are concerned.)
Perturbation Noise Levels. We study the effect of perturbation noise levels and how low the value could be in order to construct notable harm. As shown in the left three subfigures in Figure 2, under different setups, the attacks are highly successful even at a low level of noise (e.g., $\epsilon=0.05$ or even lower). We can also see the rise in ECE is sharp when $\epsilon$ increases, and it plateaus quickly.
Iterations. The numbers of iterations can help measure the efficiency and cost of attacks (in terms of both time and computing expenditure). The last subfigure in Figure 2 shows ECE vs. iteration numbers when applying MMA to the victim models on CIFAR-100. In Appendix G.3, we provide more detailed comparison, showing that the calibration attacks consistently produced a higher degree of miscalibration compared to the original unaltered SA. ECE begins to saturate at 500 iterations, but even at 100 iterations the victim models become heavily miscalibrated.
@@
4.5 Qualitative Analysis
GradCAM Visualization.
In Figure 3 we show the coarse localization maps produced with GradCAM (Selvaraju et al., 2017), which highlights the most important regions that a model relies on to make prediction using the gradients from different layers of a network. We apply GradCAM to our ResNet-50 models that are fine-tuned and tested on Caltech-101, based on the standard attack settings discussed in Section C. We choose images where the attacks led to large change in predicted confidence (at least 10%). Figure 3 shows several representative images. We can see that the coarse localization maps have minimal to no noticeable changes after the adversarial images are produced, especially in the case of the overconfidence attacked images. This analysis shows that it could be difficult to identify the attacks based on gradient visualization methods.
Figure 3: The GradCAM visualizations shows the image regions most responsible for the decisions of ResNet-50 before (top row) and after (bottom row) attacks. The left three images are under underconfidence attacks and the right three the overconfidence attack.
t-SNE.
Appendix G.4 provides detailed t-SNE visualization on the effect of different forms of calibration attacks. The visualization shows that overconfidence attack causes the representations for different classes to be split apart as much as possible, while the underconfidence attack causes a more jumbled representation with most data points falling closely to the decision boundary. The visualization shows that the attacks achieve their intended goals.
4.6 Confidence Certification.
We follow the method in Kumar et al. (2020) to compare the lower bound of expected confidence of a smoothed classifier between base images and their underconfidence and overconfidence attacked counterparts. We find that the lower bound is not significantly different between adversarially attacked images and their originals, though the accuracy of the smoothed classifier is affected. This means that methods for determining certified lower bounds are not strongly affected by calibration attacked samples (similar to conventional adversarial samples), but limitations in efficiency and accuracy of these methods means that establishing defence method against such attacks is still paramount. A detailed discussion is in Appendix G.7.
5 Defending Against Calibration Attacks
We compare a wide range of recalibration and defence methods under the setup of the MMA attacks, which, as shown above, are among the most effective calibration attacks.
Specifically, for post-calibration methods, we include Temperature Scaling (TS) (Guo et al., 2017), Multi-domain Temperature Scaling (MD-TS) (Yu et al., 2022), and calibration with splines (Splines) (Gupta et al., 2021). For training-based regularization methods we include two effective models, DCA (Liang et al., 2020) and SAM (Foret et al., 2021). Regarding adversarial defence methods, we test the top-3 SOTA models under the $l_{\infty}$ attack for CIFAR-100, using WideResNet on the RobustBench leaderboard (Croce et al., 2021), which are only available for CIFAR-10 and CIFAR-100, hence we run over CIFAR-100 to compare with our previous baselines. We further include a recent post-process defence called Adversarial Attack Against Attackers (AAA) (Chen et al., 2022) in addition to the the most common defences in the form of a PGD-based adversarial training (AT) (Madry et al., 2018). We also included the two defences proposed for calibrations attacks: CAAT and CS, which are introduced in Section 3.3. (The experiment setup is described in Appendix C, and the detailed description of these baselines are in Appendix D.2).
Results and Analyses. Table 4 shows the experiment results of query efficiency, accuracy, and the ECE and KS errors, before (PrECE and PrKS) and after the attacks (PsECE and PrKS), using different recalibration and defence models. The best performances of post-attack ECE (PsECE) and KS (PsKS) are marked in bold, and the second best are marked with underlines.
We can see that CS is overall the strongest methods at maintaining low calibration errors on post-attack datapoints. It showed the best post-attack calibration performance in five out of six setups on PsECE, and ranked among top 2 in all the other setups. The other top calibration performances are distributed among AAA, Spline, and AT. Knowing the property of calibration attacks and organizing defences accordingly is helpful to ensure a better defending result.